Sample Size: n = 5

First observation (x1) : 49

Second observation (x2) : 52

Third observation (x3) : 34

Fourth observation (x4) : 85

Fifth observation (x5) : 43

Sum ( Σx ) : 263

Sum of squares ( Σx2 ) : 15335

What is the lower quartile (Q1) of your sample data?
Answer #1

Answer

For calculating the value of lower quartile or first quartile, first we need to arrange the given data in ascending order

So, new order becomes

34, 43, 49, 52, 85

We have odd number of observations, so median will be the center

So,median = 49 (because its in the center)

Now, we have lower half = 34, 43 and upper half= 52, 85

Now, we have only two observations in lower half, so lower quartile would be the mean or average of these two observations

So, lower quartile = (34+43)/2 = (77/2) = 38.5

Thus, required lower quartile is 38.5